In how many ways can five students be seated in row of eight seat if a certain two students
a) insist on setting next to each other
b) Refuse to sit next to each other

Answer by Edwin McCravy(20054)

In how many ways can five students
Suppose the students are A,B,C,D and E
and the seats are 1,2,3,4,5,6,7 and 8

be seated in row of eight seat if a certain two students
a) insist on setting next to each other
Suppose A and B insist on sitting together.

Case 1:  A sits left of B
There are 7 ways to choose their seats: 1&2,2&3,3&4,4&5,5&6,6&7, and 7&8
That's 7 ways:
Case 2:  A sits right of B
That's also 7 ways.

That's 14 ways to place A and B together

There are now 6 seats left and 3 students C,D, and E left to seat.

For each of those 14 ways to seat A and B, there are
6 ways to seat C.  That's 14*6 ways to seat A,B and C.

There are now 5 seats left and 2 students D and E left to seat.

For each of those 14*6 ways to seat A,B and C, there are
5 ways to seat D.  That's 14*6*5 ways to seat A,B,C and D.
 
There are 4 seats left and 1 student, E, left to seat.

For each of those 14*6*5 ways to seat A,B,C and C, there are
4 ways to seat D.  That's 14*6*5*4 ways to seat A,B,C,D and E.

Answer: 14*6*5*4 = 1680 ways.

b) Refuse to sit next to each other
There are 8 ways to seat A, 7 ways to seat B, 6 ways to seat C,
5 ways to seat D and 4 ways to seat E.

That's 8*7*6*5*4 = 6720 ways anybody can sit in any seat.

We subtract 1680 ways A and B sit together.

Answer 6720-1680 = 5040.

Edwin